### v1

# Assessing Language Comprehension in Large Language Models Using Construction Grammar

Abstract
Large Language Models, despite their significant capabilities, are known to fail in surprising and unpredictable ways. Evaluating their true ‘understanding’ of language is particularly challenging due to the extensive web-scale data they are trained on. Therefore, we construct an evaluation to systematically assess natural language understanding (NLU) in LLMs by leveraging Construction Grammar (CxG), which provides insights into the meaning captured by linguistic elements known as constructions (Cxns). CxG is well-suited for this purpose because provides a theoretical basis to construct targeted evaluation sets. These datasets are carefully constructed to include examples which are unlikely to appear in pre-training data, yet intuitive and easy for humans to understand, enabling a more targeted and reliable assessment. Our experiments focus on downstream natural language inference and reasoning tasks by comparing LLMs’ understanding of the underlying meanings communicated through 8 unique Cxns with that of humans. The results show that while LLMs demonstrate some knowledge of constructional information, even the latest models including GPT-o1 struggle with abstract meanings conveyed by these Cxns, as demonstrated in cases where test sentences are dissimilar to their pre-training data. We argue that such cases provide a more accurate test of true language understanding, highlighting key limitations in LLMs’ semantic capabilities. We make our novel dataset and associated experimental data including prompts and model responses publicly available.
Assessing Language Comprehension in Large Language Models Using Construction Grammar
Wesley Scivetti Georgetown University wss37@georgetown.edu Melissa Torgbi University of Bath Austin Blodgett Army Research Lab
Molly Shichman Army Research Lab Taylor Hudson Army Research Lab Claire Bonial Army Research Lab Harish Tayyar Madabushi University of Bath
1 Introduction
Understanding the language faculties of Large Language Models (LLMs) has become increasingly important with their introduction into the mainstream. LLMs have made surprising performance gains on a wide swath of NLP tasks, but continue to fail in seemingly unpredictable and idiosyncratic ways. Linguistic theory is a possible lens for understanding these failures and LLMs more generally. Indeed, analysis of transformer language models has shown that they implicitly encode substantial linguistic information (e.g., Rogers et al., 2021; Manning et al., 2020). However, there is considerable doubt about the compatibility of traditional generative linguistic theories with LLMs (Piantadosi, 2023).
This work tests model ‘understanding’ through the creation of specific evaluation tasks based in CxG (Goldberg, 1995; Croft, 2001). CxG encompasses a family of linguistic theories which posit that morphemes, words, idioms, and even schematic structures in language can be represented as simple form-meaning pairs, or constructions (Cxns), at different levels of granularity and abstraction. Because Cxns denote semantics associated with clear syntactic patterns, they provide a mechanism to develop test instances of true ‘understanding.’ Specifically, we create an NLI task where understanding the Cxn is necessary for the model to complete the task. For example, consider the following instance of the let-alone Cxn, which invokes a scale of a property (here ease of consistency), where the slot preceding “let alone” is lower on that scale than the slot after: “It is difficult enough for an individual to be consistent let alone a society.” To test if a model understands this Cxn, we convert this to an NLI problem by using the original sentence as the premise and introducing the hypothesis “It is easier for a society to be consistent than an individual.” The NLI task is then to determine if the hypothesis follows from the premise (entailment), contradicts it (contradiction), as in the case above, or if no conclusion can be drawn (neutral). Performing this task requires both knowledge of a Cxn’s meaning and reasoning over that meaning and the surrounding context.
The use of Cxns to evaluate LLM ‘understanding’ of language offers two significant advantages. First, it allows for a systematic evaluation of language understanding by leveraging the well-studied semantic characteristics of the Cxn itself, such as the contrastive scale set up by let-alone illustrated above. Second, it enables the systematic collection of test cases that are easily understood by native speakers due to familiarity with the underlying Cxn, but are relatively uncommon or do not capture statistical co-occurrence. This second advantage is particularly important, as modern LLMs are trained on web-scale corpora, making it challenging to evaluate their performance on cases they have not been previously exposed to.
1.1 Contributions
This work aims to explore the extent to which LLMs truly ‘understand’ language through the lens of CxG, specifically targeting understanding of a variety of 8 English phrasal Cxns. This study differs from previous research in several significant ways that are highlighted by the following contributions:
1.
We create multiple novel datasets using constructional information designed to evaluate the capabilities of LLMs.
2.
Rather than focusing on the metalinguistic task of identifying Cxns, as most prior works have done, we use downstream reasoning and the well established natural language inference (NLI) task to evaluate LLM ‘understanding’ of the underlying meaning communicated by a Cxn.
3.
We test reasoning in challenge scenarios where memorization and pattern matching from large pre-training datasets can provide less assistance.
4.
We test on the largest models currently available, including GPT-4o.
The structure of this paper is as follows: We begin with a brief introduction to CxG (§2.1) and other relevant literature (§2.2). We then introduce our experimental design and its grounding in CxG theory (§3). We detail each of three different construction-based experiments (§4, §5, and §6), along with error analysis (§7). Finally, we provide an overview of our results and their implications (§8), before concluding with possible avenues of future work (§9). Overall, our evaluations indicate that LLMs struggle to understand all the nuances of constructional meaning, particularly in situations where they must generalize meaning to examples that are not statistically likely.
2 Related Work
2.1 Construction Grammars
Although there are several somewhat differing schools of CxG, most emphasize the usage-based nature of language acquisition as well as the fundamental unit of language: the Cxn. Cxns are pairings of form and meaning at all levels of language; thus, there are morphological Cxns, lexical Cxns, and phrasal Cxns. In this paper, we focus on phrasal Cxns. In the CxG account, speakers use domain-general cognitive processes to incrementally generalize over frequent Cxns, beginning with single-word Cxns, to novel usages, arriving at the ability to interpret even rare and previously unseen instances of a Cxn. Therefore, CxG provides an effective method of evaluating the true generalization capabilities of LLMs. We provide a more detailed description of CxG tenets in Appendix A.
| Cxn Name | Example |
| Causative-With | Freshly ground coffee beans filled the room with a seductive, earthy aroma. |
| Caused-Motion | But we also exported nickel to the United States. |
| Comparative-Correlative | The more I studied, the less I understood. |
| Conative | Jake sipped at the jug and didn’t answer. |
| Intransitive Motion | Armed troops marched to the substations and turned the power back on. |
| Let-Alone | None of these arguments is notably strong, let alone conclusive. |
| Resultative | He hammered the metal flat. |
| Way-Manner | A middle-aged man eased his way into the room. |
2.2 CxGs and LLMs
Starting with CxGBert (Tayyar Madabushi et al., 2020), there has been substantial past work on probing language models’ understanding of Cxns.111These works typically take a set of Cxns from CxG theoretical literature for their experiments. See Dunn (2017, 2023) for related work which attempts to discover Cxns from embedding spaces using a bottom-up approach. These works have typically focused on either a single Cxn or a handful of Cxns, like aann (Mahowald, 2023; Chronis et al., 2023), comparative-correlative (Weissweiler et al., 2022), and more schematic phrasal Cxns (Li et al., 2022; Veenboer and Bloem, 2023). Tseng et al. (2022) focus on Cxns in Taiwanese Mandarin, a notable exception to the works above which focus on English. Zhou et al. (2024) show that LLMs fail to adequately distinguish examples of the causal-excess Cxn from similar patterns, and introduce NLI as a proxy task for understanding constructions, though their results are limited to the Causal-Excess construction. Bonial and Tayyar Madabushi (2024) provide a corpus of 10 Cxns across varying levels of schematicity, showing that GPT-4 is not proficient at grouping sentences according to constructional similarity. Of particular relevance to our work is Weissweiler et al. (2022), who define a set of bias and calibration tests for the comparative-correlative, which we adapt as challenge reasoning tasks for the comparative-correlative, let-alone, and way-manner Cxns.
3 Theory-Informed Experimental Design
CxG’s usage-based account of the development of a grammar is particularly relevant to the evaluation of LLMs because language models learn through memorization, frequency and collocational properties. However, it is not clear the extent to which LLMs can recognize and generalize over the formal properties of language in order to ascribe shared semantics to novel usages (e.g., resultative: “The jackhammer pounded us deaf”). Indeed, such an ability to generalise would be a key indicator of ‘understanding.’ To explore this question, we leverage the CoGS corpus Bonial and Tayyar Madabushi (2024), which is a collection of about 500 corpus instances, roughly balanced across 10 different Cxns.
CoGs consists a limited set of carefully curated Cxns chosen for their broad coverage of the English language. These Cxns collectively represent a significant portion of English usage and provide an effective basis for evaluating LLMs in both common and uncommon linguistic scenarios. Our primary goal is not to test LLMs on specific Cxns, but rather to use constructional information to design targeted test cases that are easy for humans to understand. This allows a more precise evaluation of LLMs, resulting in the number of Cxns we use being less critical to allowing us to generalize our findings to the English language as a whole. Of the Cxns in CoGS, we construct our datasets with 8 of the 10 Cxns, shown in Table 1. In designing our dataset, we faced a trade-off between quality and quantity. While generating a large number of test samples semi-automatically was an option, we prioritized creating a smaller, carefully curated set of high-quality examples. This not only allowed for meticulous manual verification of the curated data but also allowed us to manually validate the results, which would be infeasible with a larger dataset. By minimizing biases and errors in both dataset creation and evaluation we ensure the reliability of our findings, even if it comes at the cost of a reduced number of examples.
4 Experiment 1: Constructional NLI
4.1 Datasets
We take Natural Language Inference (NLI) as a good starting point for investigating LLM capabilities regarding the semantics of Cxns, following (Zhou et al., 2024). NLI tasks present a premise, taken to be true, and a hypothesis, which can be entailed by the premise, contradicted by the premise, or neutral (neither contradicted nor entailed). Thus, by targeting entailment and contradiction, NLI effectively probes for a functional understanding of a text. NLI is also a widely accepted task for testing NLU in the NLP community, as evidenced by the many shared evaluations ranging from the general evaluation of distributional semantic models Marelli et al. (2014) to the evaluation of LLMs for biomedical applications Lee et al. (2024).
In particular, we introduce an NLI dataset which specifically probes the semantics of our Cxns. We do this by templatically generating hypotheses which make use of certain slots within a Cxn, and verifying the resultant sentences manually to ensure quality. These template-based hypotheses have predictable relations to the premises. For example, consider the premise: "He hammered the metal flat". This is a resultative Cxn, which has the meaning of an action which causes a change in state. In this case, the action verb hammered leads to the metal to have a resulting state of flat. Regarding the syntactic form of a phrasal Cxn, we can use a constructional template to define the slots that are filled by a given Cxn. A general template for the resultative is shown in Example 4.1, and its application to the above sentence in Example 4.1.
. [SBJ1 [V2 OBJ3 ADJ4 ]VP]5
. [He1 [hammered2 the metal3 flat4 ]VP]5
Given constructional examples like those in Example 4.1, our goal is to produce NLI tuples that consistently target the construction’s meaning. We do this by manipulating the slots in the Cxn templatically to produce hypotheses with consistent relations to the premise. For instance, consider the hypothesis "The hammering did not cause the metal to become flat." This is clearly a contradiction to the above premise and is directly in conflict with the meaning of the Cxn. We can produce this hypothesis, and similar hypotheses, by following the template in Example 4.1. We can see this template reflecting the above hypothesis in Example 4.1.
. [The [V]2-ing did not cause [OBJ]3 to become [ADJ]4].
. [The [hammer]2-ing did not cause [the metal]3 to become [flat]4].
For each Cxn, we create a template to generate NLI hypotheses that target constructional meaning.222There are many ways to create NLI hypotheses beyond the templates we utilize for our hypotheses. We also create a more freeform NLI dataset given our constructional premises, finding similar results. These templates produce hypotheses that have predictable relations, and are directly tied to the semantics of the Cxn. See Appendix B for details. See Appendix E for the NLI templates for each Cxn, along with examples. Overall, our process for creating the constructional NLI dataset can be summarized as:
1.
Extract natural constructional examples from CoGS Bonial and Tayyar Madabushi (2024).
2.
Create general templates for NLI hypotheses for each Cxn.
3.
Generate hypotheses for each example Cxn given the corresponding templates.
4.
Manually validate the resulting dataset.
All templatically generated hypotheses and relations were validated by an author. Once the dataset was created, a second author annotated the relations, achieving an Inter-Annotator Agreement (IAA) of 90% agreement. Descriptive statistics for this dataset along with all other datasets can be found in Appendix C.
4.2 Empirical Evaluation and Analysis
We test three OpenAI models333https://platform.openai.com/docs/models on our constructional NLI dataset. We test the GPT-4o-2024-05-13 and GPT-3.5-turbo-0125 models as well as the recently released o1-preview-2024-09-12 model. These models were chosen due to their large sizes making them illustrative examples of the capabilities of state-of-the-art LLMs in general. We also test two LLaMa models, LLaMA-3-8b-instruct and LLaMA-3-70b-instruct444Full results for the LLaMa experiments can be found in Appendix J. We test 3 main scenarios: zero-shot, in-context learning with examples randomly selected from Stanford NLI (SNLI, Bowman et al. 2015), and in-context learning with Constructional NLI (CxNLI, our dataset). We additionally perform Chain-of-Thought (CoT) experiments (Wei et al., 2022b) in each of these scenarios, in an attempt to more precisely explore gaps in model capabilities. Surprisingly, we find that Chain-of-Thought does not lead to performance gains on our dataset.555Full results for CoT are reported in Appendix F A summary of our results for this experiment are reported in Table 2.666We experiment with a variety of prompt formats and report results for the best performing prompts. Details on prompt variations and results can be found in Appendix D
| Setting | IC Data | GPT Accuracy | ||
| 3.5 | 4o | o1* | ||
| Zero-shot | None | 0.6 | 0.88 | 0.90 |
| One-shot | CxNLI | 0.69 | 0.9 | - |
| Three-shot | CxNLI | 0.79 | 0.96 | 0.90 |
| One-shot | SNLI | 0.58 | 0.86 | - |
| Three-shot | SNLI | 0.59 | 0.86 | 0.89 |
Overall, we see that performance is quite high even in the zero-shot setting for GPT-4o and GPT-o1. We also observe that GPT-4o consistently perform better than GPT-3.5. Adding additional examples of Cxns for in-context learning boosts performance, while additional SNLI examples do not boost performance. This is especially true for GPT-3.5, which benefits substantially more from in-context learning from CxNLI. This reliance on in-context learning indicates that our datasets test a different axis of semantic knowledge than more general datasets like SNLI.
5 Experiment 2: Challenge Constructional NLI
5.1 Dataset
In the previous section, we show that GPT-4o and GPT-3.5 perform impressively on an NLI task that specifically targets constructional semantics. Many of these constructions are common to the English language, and the templates we generate target aspects of meaning that are highly salient for the construction. However, we wish to test LLM generalization of constructional knowledge even in settings with unexpected uses of the construction. In order to do this, we develop a more challenging and NLI dataset (CxNLI-challenge). For 5 of our 8 Cxns, we introduce more challenging examples, which share a surface syntax with our Cxns, but convey a different meaning. To demonstrate this, consider the following two examples:
. He hammered the metal flat. (resultative)
. I bought the apples fresh. (depictive)
In the above two examples, it is clear that their syntactic forms appear identical. Both have a subject pronoun, a verb, an object noun phrase followed by an adjective. However, the two Cxns convey different meanings: In Example 5.1 the adjective is the result of the action of the verb, whereas in Example 5.1, the adjective is the state of the noun during the action of the verb, but it is not the resulting state of the action. This difference in meaning is associated with two different Cxns, specifically the resultative and the depictive. We can tease apart this difference in meaning by leveraging our template based hypotheses from our CxNLI dataset in 4. Consider the following templatic hypotheses:
. [The [hammer]2-ing caused [the metal]3 to become [flat]4]. (entailment)
. [My [buying]2-ing caused [the apples]3 to become [fresh]4]. (contradiction)
As we can see in Examples 5.1 and 5.1, templatically generating hypotheses for these two examples leads to different relations to the premises. We create this more challenging dataset (CxNLI-Challenge) by gathering "false positives" of Cxns from corpora based on simple pattern matching queries. These "false positive" constructions are those which on the surface appear to be the same as the constructions which we are testing, but have drastically different meanings. We templatically generate hypotheses for these false positives, which produces a dataset very similar in form to CxNLI (described in 4), but with different relations between premises and hypotheses, due to the different meanings of the false positive constructions. In many cases, the resulting hypotheses are rarer phrasings for the challenge Cxns than they are for their intended Cxns from CxNLI. We hypothesize that this feature will make them more difficult for the model, as they are less likely to be highly represented in the pretraining data for the LLM. For humans, interpretation is possible because we are able to generalize the meanings coming from certain slots in the Cxn even in unusual contexts. After the final version of this dataset was created, a second author evaluated the dataset, achieving an IAA of 83% with the original judgements.
5.2 Empirical Evaluation and Analysis
We utilize the CxNLI-challenge dataset in two settings: as in-context examples for our CxNLI dataset from Section 4, and as an evaluation dataset. We first examine what effect, if any, these challenge examples have on CxNLI performance when they are used as the in-context learning examples for our previous dataset. As we show in Table 3, for GPT-4o, performance using CxNLI-challenge as in-context learning for our CxNLI does not boost or harm performance significantly, and is roughly equivalent to performance in the zero-shot setting. On the other hand, performance on CxNLI drops significantly for GPT-3.5 when it is prompted with CxNLI-challenge. This indicates that GPT-3.5 does not have the same constructional knowledge as GPT-4o, and is easily misled by surface syntactic similarity.777We show for best performing LLM prompts in Table 3. See Appendix D for results for full range of prompts for all experiments. We also experiment with CoT for this dataset, finding that it does not substantially affect performance.
| Setting | IC Data | GPT Accuracy | |
| 3.5 | 4o | ||
| Zero-shot | None | 0.60 | 0.88 |
| One-shot | CxNLI-challenge | 0.38 | 0.89 |
| One-shot | CxNLI-challenge,CxNLI | 0.7 | 0.9 |
We further test if LLMs can perform NLI successfully on these challenge examples themselves. As we can see in Table 4, performance is significantly lower on our challenge NLI dataset in almost every prompt setting. The difference in performance is stark: while these examples are admittedly more unnatural sounding and indeed more difficult for humans, the ceiling of human performance (IAA 83%) is well above current LLM performance, even for GPT-4o and GPT-o1. Again we see a large difference between GPT-4o and GPT-3.5, with GPT-3.5 performing below chance in all task settings. We also see that GPT-o1 performs worse than GPT-4o on this task.
| Setting | IC Data | GPT Accuracy | ||
| 3.5 | 4o | o1 | ||
| Zero-shot | None | 0.26 | 0.58 | 0.46 |
| One-shot | CxNLI | 0.26 | 0.53 | - |
| Three-shot | CxNLI | 0.31 | 0.57 | 0.45 |
| One-shot | SNLI | 0.3 | 0.55 | - |
| Three-shot | SNLI | 0.28 | 0.53 | 0.46 |
6 Experiment 3: Challenge Constructional Reasoning
6.1 Dataset
The above experiment on challenge NLI is only possible for some Cxns. For 5 of our 8 Cxns, we successfully create challenge NLI examples based on shared syntactic forms with other Cxns. However, for 3 of our 8 Cxns (let-alone, comparative-correlative, way-manner), creating a challenge dataset is not straightforwardly possible using form-based “false positives.” A clear example of this is the let-alone Cxn: its syntactic pattern is unique to the Cxn, and there are no examples of this exact syntactic pattern that aren’t instances of the let-alone Cxn. Since we cannot create challenge NLI examples for these 3 Cxns with templates, we must find another method for further investigating the semantic understanding of these Cxns.
In order to tackle this problem, we draw on the semantic experiments from Weissweiler et al. (2022). They lay out a group of 7 tests which are designed to probe the semantic understanding of the comparative-correlative while accounting for a range of potential form-based biases and confounding factors. They find that a set of BERT-style models fail at most of these tests. We extend the tests outlined in Weissweiler et al. (2022) to naturally occurring instances of the comparative-correlative, as opposed to artificial examples used in previous work, and we expand our test set to include similar tests for the let-alone and way-manner Cxns. We also change the task to a Question-Answering (QA) task, as opposed to a Masked Language Modeling (MLM) task to better suit the LLMs we test. In our QA scenario, there are always 3 choices: 2 are based on opposing interpretations of the Cxn, and the third choice is “not enough information,” which the model should select when the Cxn does not provide enough information to answer the question.888Descriptions and examples of every test can be found in Appendix G
While these tests are more specific than the NLI examples above, we believe that this dataset serves a similar purpose to the challenge NLI dataset: to more rigorously probe the meaning of Cxns while controlling for the form as much as possible. After creating this dataset for the let-alone, way-manner and comparative-correlative Cxns, we had an additional author annotate the dataset, achieving IAA of 78%.
6.2 Empirical Evaluation and Analysis
| Cxn | GPT Accuracy | ||
| 3.5 | 4o | o1* | |
| Comparative-Correlative | 0.47 | 0.70 | 0.61 |
| Let-Alone | 0.56 | 0.58 | 0.65 |
| Way-Manner | 0.54 | 0.67 | 0.67 |
[…]
Somewhat surprisingly, we find that LLMs remain unable to solve the tasks of Weissweiler et al. (2022) at high accuracy, even in the Chain-of-Thought setting.999More details on CoT experiments in Appendix F. We find that this low performance is consistent across most of the tests from Weissweiler et al. (2022), although there is wide variation on specific test performance depending on the prompt. Unsurprisingly, we again find that GPT-4o and GPT-o1 outperforms GPT-3.5, though the difference is not as stark as in prior experiments.
7 Error Analysis
7.1 Constructional NLI and challenge NLI
In §4, we describe our Constructional NLI experiments, which find that GPT-4o and GPT-o1 are extremely proficient at Constructional NLI (CxNLI) while GPT-3.5 lags behind substantially. In §5, we find a similar trend for challenge NLI (CxNLI-challenge), but show that even GPT-4o and GPT-o1 do not perform well on CxNLI-challenge examples. For example, though its performance on the CxNLI resultative is near perfect, it struggles with CxNLI-challenge false positives for the resultative in some cases, as shown below.
[…]
Hypothesis: The apples were completely fresh before I bought them.
Correct Response: Entailment
Model Response: Contradiction
Here, we investigate if some Cxns are harder for LLMs than others. In Figure 1 we report the accuracy by Cxn in our CxNLI and CxNLI-challenge datasets.101010Accuracies are from the highest performing prompt. In general, the same Cxns are difficult across the CxNLI and CxNLI-challenge datasets. Overall, we see that let-alone and comparative-correlative are the weakest Cxns for GPT-4o, though it is strong across the board with a minimum accuracy of 88%. We show an example of GPT-4o misunderstanding the scale of let-alone in Example 10.
[…]
Hypothesis: Beecher’s reputation as a Man of God was easier to accept than his reputation as a preacher.
Correct Response: Contradiction
Model Response: Entailment
GPT-3.5 is much more variable by Cxn, with a maximum accuracy of 92% for the conative and a minimum of 67% for let-alone.
7.2 Challenge Constructional Reasoning
In §6, we show that GPT-4o, GPT-o1 and GPT-3.5 struggle on a QA formulation of a broad set of challenge Cxn reasoning tasks. After examining the errors across all tests, we highlight that there is significant variation in test-by-test performance based on the input prompt. We highlight two possible prompts here, Prompt 1 and Prompt 4.111111Description of prompt variations and results for all experiments can be found in Appendix D The key difference in these prompts is how they task the model with uncertainty. Since some of the QA tests are designed to be unanswerable given the context, the model is supposed to choose the answer not enough information (always choice C in our QA setup). However, we find that the model is unable to balance between two extremes. In some cases, it answers “not enough information” when interpreting the Cxn is sufficient for providing a likely answer, and in other cases, it makes incorrect assumptions about the relationship between elements of the Cxn and context. We can see this in Figure 2. Prompt 4 includes the line “If there is not enough information, choose C,” and does far better on the tests where there is in fact not enough information, but substantially worse on the tests where the Cxn is enough information to provide an answer. An example of this failure is show in Example 11, where the model misinterprets a nonce adjective as somehow relating to the scale of the Cxn, when in fact there is not enough information to answer the question. As a whole, these results show that in our current experimental setup, we cannot get the models to focus specifically on meanings associated with Cxns as opposed to broader context and world knowledge. This leads to incorrect reasoning in cases where the close interpretation of the Cxn outweighs the other factors needed to determine the correct answer.
[…]
Correct Response: Not enough information
Model Response: Pay More
8 Discussion
Overall, these experiments show the limits of LLM functional understanding of commonplace English Cxns. Regarding our constructional NLI experiments, our results consistently show that LLMs perform better when they encounter in-context examples similar to the actual task they are required to perform, as opposed to to typical SNLI examples. In fact, an increase in the number of SNLI examples leads to deteriorating performance. This result initially seems counterintuitive, as both types of examples provide the same information about the underlying NLI task. However, this can be explained by viewing LLMs’ use of examples as akin to fine-tuning Dai et al. (2023), and instruction tuning—which aims to help models understand instructions Wei et al. (2022a)—as facilitating them to learn in-context directly on instructions Lu et al. (2024). From this perspective, it becomes clear why examples with the same statistical distribution as the task at hand perform better, despite providing the same task-related information. Regarding Cxn reasoning, we find that prompt variation contributes significantly to overall reasoning over Cxns, though all prompt formulations lead to performance substantially below human levels. Overall, our findings indicate GPT-4o and GPT-o1 have substantially more awareness of Cxns than GPT-3.5 has, which may have implications for the usefulness of these respective models for linguistic research on CxG. However, we find that both GPT-4o and GPT-o1 fail to generalize their constructional knowledge sufficiently to tackle challenge scenarios, on both Constructional Reasoning and NLI challenge tasks.
9 Conclusions and Future Work
In this work, we have shown where even the latest GPT-4o and GPT-o1 models do not demonstrate functional understanding of Cxns. While GPT-4o and GPT-o1 do perform quite impressively on our constructional NLI task, they fail at both challenge NLI and challenge reasoning scenarios, which require generalization of the constructional semantics to infrequent instantiations of the Cxn. Also, we see that GPT-4o substantially outperforms GPT-3.5 in all settings. GPT-3.5 is especially susceptible to being misled by shallow form-based patterns as opposed to meanings, and in-context learning is especially crucial for GPT-3.5. Overall, these experiments show that the constructional awareness of both GPT-4o and GPT-o1 are far more robust than that of GPT-3.5, but their ability to generalize the meaning of Cxns to new contexts, and thus its “understanding”, still lags substantially behind that of humans.
Thus, our targeted series of experiments demonstrated that LLMs do process constructional semantics up to a point, yet our challenge datasets revealed the breaking point of understanding—where speakers are able to generalize the appropriate semantics to constructional slots filled by pragmatically atypical lexical items, but LLMs are much less proficient at this generalization. Overall, we find that CxG serves as a valuable theoretical lens for probing the functional language understanding of LLMs and comparing this with human linguistic knowledge. Greater contributions to resources such as corpora of Cxns will facilitate empirical data on which constructional understanding can be evaluated with more detail.
Limitations
This work is limited in that we only evaluate our methods on English. More work is needed on the targeted evaluation of LLM performance using Cxn information in non-English settings. Our tasks are only one possible method for investigating LLM understanding of Cxns. Expanding research to include complementary methodologies will be necessary to build a complete picture of LLM knowledge in relation to CxG. This work can also be extended beyond the 8 Cxns that we use to generate our dataset, although these were selected for the extensive coverage of the English language. Also, while we consciously choose to create a smaller, more carefully curated dataset that also allows for careful expert manual evaluation, there is scope to increase the size of our dataset, which we leave to future work.
Ethics
LLMs are extremely expensive to train and run. The compute costs associated with LLMs have a nontrivial environmental impact which should not be ignored. Furthermore, due to their large-scale training data, they can reflect and propagate harmful social biases in their responses if they are not properly aligned and moderated. Furthermore, there is risk of LLMs having a negative societal impact if their widespread deployment is done without proper consideration for the lives of people. While there are risks in the use and proliferation of LLMs in general, we do not believe this work incurs any specific additional risks. Despite the overall risks and dangers, we believe this research is worthwhile in order to better understand the language systems of LLMs and compare and contrast LLM language understanding with that of humans. We honor the code of ethics.
Appendix A Construction Grammers
CxG has particular explanatory power with respect to phrasal constructions, such as the resultative construction: “The jackhammer pounded us deaf.” Generative linguistic theories (e.g., Chomsky (2014)) would generally analyze a transitive sentence with one verbal head (“pounded”) that licenses the arguments of the sentence. “Pounded” generally licenses an agent subject (here, “jackhammer”) and potentially a patient direct object. Generative approaches argue that this information about the verb is memorized and stored in the lexicon, while combinatory rules of how to put lexical items together are stored in a separate syntax module of language processing. However, unless a special grammatical rule or sense of the verb is postulated, there is nothing to explain why “us” is not the thing pounded here, or what licenses the adjective “deaf.” Nonetheless, native speakers have no problem recognizing the special formal and semantic properties of this Cxn, namely that it entails a pounding event that causes a change in state of “us” resulting in the state of “deaf.”
In contrast to Generative linguistic theories (e.g., Chomsky (2014)), CxG posits that speakers acquire and store constructions, which notably account for not only the semantic properties of the unit but also the formal syntactic properties. Constructions at all levels of language are learned through language usage; thus, in lieu of grammatical rules accounting for the grammaticality of a particular structure, frequency plays an important role in what ‘sounds right’ or is grammatical to a speaker. In the CxG account, speakers acquire single-word constructions that they are frequently exposed to (e.g., “milk”), but then generalize from that to recognize how an acquired holophrastic Cxn falls into slots of larger, more complex constructions (e.g., “want milk,” “baby want milk,” “want up,” “I want to go”). In this process of generalization, speakers build up a taxonomically organized set of the related constructions of their language, or a constructicon. Speakers use domain-general cognitive processes to incrementally generalize over such frequent constructions to novel usages, arriving at the ability to interpret even rare and previously unseen instances of a Cxn.
Appendix B Templatic Constructional NLI vs Freeform NLI
| Setting | IC Data | Dataset | Accuracy | |
| GPT-3.5 | GPT-4o | |||
| Zero-shot | None | Templatic | 0.6 | 0.88 |
| One-shot | CxNLI | Templatic | 0.69 | 0.9 |
| Three-shot | CxNLI | Templatic | 0.79 | 0.96 |
| One-shot | SNLI | Templatic | 0.58 | 0.86 |
| Three-shot | SNLI | Templatic | 0.59 | 0.86 |
[…]
| One-shot | CxNLI | Freeform | 0.78 | 0.91 |
| Three-shot | CxNLI | Freeform | 0.83 | 0.94 |
| One-shot | SNLI | Freeform | 0.7 | 0.89 |
| Three-shot | SNLI | Freeform | 0.69 | 0.9 |
In addition to our templatically generated constructional NLI dataset, we construct a more "freeform" NLI dataset by training a person to create hypotheses that are intended to be more similar in form to hypotheses from SNLI. Both our "freeform" and templatic datasets are constructed with real-world corpus data of constructions, primarily coming from the COGS dataset (Bonial and Tayyar Madabushi, 2024), with supplementary data coming from Corpus of Contemporary American English (COCA, Davies 2010), and the English Corpus from the Web, or EnCOW (Schäfer and Bildhauer, 2012; Schäfer, 2015). Within both of these datasets, each premise includes one of 8 total constructions: the comparative-correlative Cxn, the let-alone construction, the way-manner construction, the causative-with construction, the conative construction, the resultative construction, the caused motion construction, and the intransitive motion Cxn. We include a roughly balanced sample of each construction, with all premises taken from corpus data. These constructions cover a wide range of schematicity meaning that they have different levels of lexicalization/abstractness.
| Cxn Name | Cxn Template | NLI Hypothesis Template |
| Causative-With | [SBJ1 [V2 OBJ3 with-PP4 ]VP]5 | SBJ1 did not cause OBJ3 to contain OBJ-of-PP4 |
| Caused-Motion | [SBJ1 [V2 OBJ3 PP4 ]VP]5 | OBJ3 did not change locations. |
| Comparative-Correlative | [[the1 [Comparative Phrase]2 REST-CLAUSE3]C1 [[the4 [Comparative Phrase]5 REST-CLAUSE6]C2]7 | The amount [Comparative Phrase]2 is positively/negatively correlated with the amount [Comparative Phrase]5 |
| Conative | [SBJ1 [V2 at-PP3 ]VP]4 | OBJ-of-PP3 was not the target of the V2-ing motion. |
| Intransitive-Motion | [SBJ1 [V2 PP3 ]VP]4 | SBJ1 V2 in a static location. |
| Let-Alone | XP1 CONJ2-3 XP4-5 | If XP1 then not XP4-5 |
| Resultative | [SBJ1 [V2 OBJ3 AP4 ]VP]5 | The V2-ing did not cause OBJ3 to become AP4. |
| Way-Manner | [SBJ1 [V2 [PRON3=1] way4 ]OBJ5 PP6]VP]7 | SBJ1 traveled PP6 without V2-ing. |
| Dataset | Ex. | Tokens | IAA |
| Cxn NLI (Freeform) | 435 | 15,401 | 92% |
| Cxn NLI (Templatic) | 435 | 15,144 | 90% |
| challenge NLI | 99 | 3,202 | 82% |
| challenge Reasoning | 740 | 41,835 | 78% |
| Prompt Type | Accuracy | |
| GPT-3.5 | GPT-4o | |
| Variation 1 | 0.74 | 0.95 |
| Variation 2 | 0.79 | 0.92 |
| Variation 3 | 0.76 | 0.94 |
| Variation 4 | 0.74 | 0.93 |
[…]
| Best Variation Rephrase 2 | 0.67 | 0.95 |
| Best Variation Rephrase 3 | 0.69 | 0.96 |
| Prompt Type | Accuracy | |
| GPT-3.5 | GPT-4o | |
| Variation 1 | 0.24 | 0.53 |
| Variation 2 | 0.28 | 0.46 |
| Variation 3 | 0.31 | 0.57 |
| Variation 4 | 0.26 | 0.54 |
| Best Variation Rephrase 1 | 0.31 | 0.57 |
| Best Variation Rephrase 2 | 0.26 | 0.55 |
| Best Variation Rephrase 3 | 0.23 | 0.54 |
| Prompt Type | Accuracy | |
| GPT-3.5 | GPT-4o | |
| Variation 1 | 0.41 | 0.61 |
| Variation 2 | 0.38 | 0.61 |
| Variation 3 | 0.45 | 0.64 |
| Variation 4 | 0.42 | 0.68 |
[…]
| Best Variation Rephrase 2 | 0.58 | 0.68 |
| Best Variation Rephrase 3 | 0.52 | 0.68 |
| Cxn Name | Cxn NLI (Freeform) | Cxn NLI (Templatic) | |
| Causative-With | Premise | Freshly ground coffee beans filled the room with a seductive, earthy aroma. | Freshly ground coffee beans filled the room with a seductive, earthy aroma. |
| Hypothesis | Stale whole coffee beans filled the room with an unattractive, otherworldly scent. | The room did not contain a seductive, earthy aroma. | |
| Relation | Contradiction | Contradiction | |
| Caused-Motion | Premise | I threw the stone across the river. | I threw the stone across the river. |
| Hypothesis | I threw the stone up and caught it again. | I caused the stone to move across the river by throwing it. | |
| Relation | Contradiction | Entailment | |
| Comparative-Correlative | Premise | The more they work, the more I will pay them. | The more they work, the more I will pay them. |
| Hypothesis | I intend to compensate my workers based on their amount of work. | Increasing the amount they work will increase the amount I pay them. | |
| Relation | Entailment | Entailment | |
| Conative | Premise | I sipped at the Heineken. | I sipped at the Heineken. |
| Hypothesis | I spit out the Heineken. | The Heineken was not the target of my sipping. | |
| Relation | Contradiction | Contradiction | |
| Intransitive-Motion | Premise | I ran around the track. | I ran around the track. |
| Hypothesis | I walked around the track. | I ran, staying in one place. | |
| Relation | Contradiction | Contradiction | |
| Let-Alone | Premise | It’s unsurprising that such an attitude failed to produce competent screenwriters, let alone exciting ones. | It’s unsurprising that such an attitude failed to produce competent screenwriters, let alone exciting ones. |
| Hypothesis | Although such an attitude failed to produce competent screenwriters, it produced plenty of exciting screenwriters. | An attitude that produces exciting screenwriters can also produce competent ones. | |
| Relation | Contradiction | Entailment | |
| Resultative | Premise | The jackhammer pounded us deaf. | The jackhammer pounded us deaf. |
| Hypothesis | The jackhammer was easy on our ears. | We were completely deaf before the jackhammer pounded. | |
| Relation | Contradiction | Contradiction | |
| Way-Manner | Premise | I yawned my way back to the Narrow Neck. | I yawned my way back to the Narrow Neck. |
| Hypothesis | I yawned at least once on my journey back to the Narrow Neck. | I traveled back to Narrow Neck without yawning. | |
| Relation | Entailment | Contradiction |
The freeform Constructional NLI tuples are constructed in order to be similar to other NLI datasets, testing general types of semantics. These tuples do not always necessitate perfect understanding of the present Cxn in order to be correct; rather, the surrounding context is sometimes enough to provide the correct answer. On the other hand, the templatic Constructional NLI dataset is designed to specifically test the meanings associated with the present constructions. Each NLI tuple makes direct use of the meaning attached with the constructions, and thus providing the correct answer to the NLI task implies correctly associating some aspect of meaning with the Cxn. These templates also allow us to easily construct our challenge examples, which is crucial for the experiment laid out in 5. Examples of Cxn and NLI templates are provided in Table 7. For full examples of NLI triples for each of the 8 constructions evaluated, see Table 12. We test GPT-4o and GPT-3.5 on our freeform Cxn NLI dataset, and find generally similar trends to the templatic Cxn NLI dataset. A results comparison across the two NLI datasets is shown in Table 6.
Appendix C Scientific Artifacts and Descriptive Statistics for All Datasets
We use the following scientific artifacts: COCA (Davies, 2010), EnCOW (Schäfer and Bildhauer, 2012; Schäfer, 2015), the CoGS dataset (Bonial and Tayyar Madabushi, 2024), and the OpenAI API, in addition to our created datasets. The COCA corpus contains 8 genres: Academic, Blog, Fiction, Magazine, News, Spoken, TV, and Web. It is intended to capture American English, though there is no guarantee that it does not also include some other varieties. Demographic information about the creators of the texts in the corpus is not always available given its scale. EnCOW is a large scale corpus of English text from the web. As such, the demographic information that it captures is not completely clear. All datasets besides COCA and EnCOW are open-source under a Creative Commons license. The institutions of the authors have valid licenses for COCA and EnCOW, permitting their use in academic settings. We use the artifacts as intended. Overall, we construct and experiment on 4 datasets: 2 constructional NLI datasets (templatic and freeform), the challenge NLI dataset, and the challenge reasoning dataset. The sizes and IAA agreement for the final datasets is reported in Table 8. All of our datasets are exclusively in English.
Appendix D Prompt Variation Experiments
For each experiment, we choose one setting to observe the impact of prompt variations on performance.121212For NLI we choose the 3-shot CxNLI setting. For challenge reasoning we choose zero-shot with a subset of all constructions. These prompt variations were explored in two sequential stages; changing the content of the prompt and changing the phrasing of the prompt. Changing the content of the prompt involved giving more details about the task, giving fewer details about the task and experimenting with the specific instructions. After selecting the best-performing prompt content, the prompt was rephrased using an LLM and variations of the prompt with new wording were used to check the robustness of the model output by observing how small variations in the prompts affect performance. We report an example of each of our prompts for variations in content in Tables 13 and 14. Results for our prompt variation experiments are shared in Tables 9, 10 and 11. Annotators were shown the exact same prompts as models when doing their annotations.
| Prompt Type | Prompt |
| Variation 1 | You are the world’s best annotator. Your task is to read sentences from a dataset, presented as the Premise in a set of triples for the Natural Language Inference (NLI) task. Also known as Recognizing Textual Entailment (RTE), NLI involves determining the inference relation between two short, ordered texts: entailment, contradiction, or neutral. Next, you will identify the Relation between the Premise and the Hypothesis, which indicates the type of entailment between the two sentences. We use numerical coding, also listed in your annotation spreadsheet as a reminder: |
| 0 – entailment – The hypothesis must be true given the premise | |
| 1 – neutral – The hypothesis may or may not be true given the premise | |
| 2 – contradiction – The hypothesis must not be true given the premise | |
| Output a single numerical value between 0, 1, or 2, corresponding to the associated relation. Output a single number only and nothing else. | |
| Variation 2 | You are the world’s best annotator. You are tasked with annotating a triple for Natural Language Inference. You must determine the inference relation between the Premise and the Hypothesis by selecting one of three numerical codes that reflect the relationship: |
| 0 – Entailment: The Hypothesis is definitely true given the Premise. | |
| 1 – Neutral: The Hypothesis may or may not be true given the Premise. | |
| 2 – Contradiction: The Hypothesis cannot be true given the Premise. | |
| Output a single numerical value between 0 and 2 inclusive, corresponding to the associated relation. | |
| Variation 3 | You are the best at understanding language inference based on Cxn grammar. You are tasked with annotating a triple for Natural Language Inference. You must determine the inference relation between the premise and the hypothesis by selecting one of three numerical codes that reflect the relationship: |
| 0 – entailment – The hypothesis must be true given the premise | |
| 1 – neutral – The hypothesis may or may not be true given the premise | |
| 2 – contradiction – The hypothesis must not be true given the premise | |
| Output a single numerical value between 0, 1, or 2, corresponding to the associated relation. Output a single number only and nothing else. | |
| Variation 4 | You are the world’s best annotator. Your task is to read sentences from a dataset, provided as the Premise in a set of triples for the Natural Language Inference (NLI) task. Also called Recognizing Textual Entailment (RTE), NLI requires determining the inference relation between two short, ordered texts: entailment, contradiction, or neutral. Your next step is to identify the Relation between the Premise and the Hypothesis, specifying the type of entailment between the two sentences. We use the following numerical coding: |
| 0 – entailment – The hypothesis must be true given the premise | |
| 1 – neutral – The hypothesis may or may not be true given the premise | |
| 2 – contradiction – The hypothesis must not be true given the premise | |
| Output a single numerical value between 0 and 2 inclusive, corresponding to the associated relation. |
| Prompt Type | Prompt |
| Variation 1 | You are the world’s best annotator. Your task is to read questions from a dataset, and determine the correct answer from the options provided using only the information given in the question. Output a single character of either A, B or C and nothing else. |
| Variation 2 | Using only the information given in the question, output a single character of either A, B, or C, corresponding to the correct answer. Output a single character only and nothing else. |
| Variation 3 | You are the world’s best annotator. Your task is to read questions from a dataset, and determine the correct answer from the options provided using only the information given in the question. Output a single character of either A, B or C and nothing else. If the answer is not clear from the context provided, output the character C. |
| Variation 4 | Using only the information given in the question, output a single character of either A, B, or C, corresponding to the correct answer. Output a single character only and nothing else. If the answer is not clear from the context provided, output the character C. |
Appendix E Example NLI Tuples
In Table 12, we show examples of our constructional NLI datasets. We show challenge Cxn NLI examples in Table 15.
| CxN Name | Example NLI Triple | |
| Conative | Premise | I watch the women, their legs crossed at the ankles, try to look as if they don’t sweat at all. |
| Hypothesis | Their legs made a crossing motion towards their ankles. | |
| Relation | Contradiction | |
| Causative- with | Hypothesis | He hit the lamp with his head. |
| Premise | He caused the lamp to contain his head. | |
| Relation | Contradiction | |
| Intransitive- motion | Hypothesis | He spoke to the workers on the street corner. |
| Premise | He changed locations by speaking. | |
| Relation | Contradiction | |
| Caused- motion | Hypothesis | I introduced her to my boss. |
| Premise | She remained in the same place. | |
| Relation | Neutral | |
| Resultative | Hypothesis | A famous emperor buried China’s scholars alive with their books. |
| Premise | Burying caused the scholars to become alive. | |
| Relation | Contradiction |
Appendix F Chain-of-Thought Results for all Experiments
We replicate each of our main experiments with the addition of Chain-of-Thought (CoT) to observe changes in performance and inspect any errors in the model’s reasoning steps. For CoT thought prompting we do not provide examples of reasoning in the prompt, instead we simply ask the model to explain "step by step". The results for each experiment have been shared in tables 16, 17, 18 and 19.
| Setting | IC Data | Accuracy | |
| GPT-3.5 | GPT-4o | ||
| Zero-shot | None | 0.60 | 0.89 |
| One-shot | CxNLI | 0.66 | 0.89 |
| Three-shot | CxNLI | 0.71 | 0.92 |
| One-shot | SNLI | 0.63 | 0.89 |
| Three-shot | SNLI | 0.66 | 0.91 |
| Setting | IC Data | Accuracy | |
| GPT-3.5 | GPT-4o | ||
| Zero-shot | None | 0.62 | 0.89 |
| One-shot | Challenge | 0.6 | 0.9 |
| One-shot | Challenge, CxNLI | 0.6 | 0.9 |
| Setting | IC Data | Accuracy | |
| GPT-3.5 | GPT-4o | ||
| Zero-shot | None | 0.29 | 0.43 |
| One-shot | CxNLI | 0.29 | 0.43 |
| Three-shot | CxNLI | 0.35 | 0.46 |
| One-shot | SNLI | 0.28 | 0.43 |
| Three-shot | SNLI | 0.31 | 0.45 |
| Cxn | Accuracy | |
| GPT-3.5 | GPT-4o | |
| comparative correlative | 0.51 | 0.68 |
| let-alone | 0.51 | 0.63 |
| way manner | 0.63 | 0.64 |
| all | 0.53 | 0.66 |
Appendix G Reasoning Test Description and Examples
There are 7 bias and calibration tests outlined in Weissweiler et al. (2022) for the comparative-correlative. However, their tests are templatically generated using one specific syntactic form of the comparative-correlative. We adapt these tests to natural examples of the construction, and apply the tests to let-alone and way-manner. We also introduce a new test called "Nonce Calibration", which is similar in spirit to the "Adjective Calibration" test but using a nonce word. A brief description of all test types is below.
1.
Base - this test presents an example of a construction, along with additional context, followed by a question. The constructional meaning combines with the context to provide the answer to the question.
2.
Recency Bias - this test reorders the Cxn and the contextual information, placing the contextual information first. The answer should be the same as Base.
3.
Vocabulary Bias - this test reverses the semantics of the scalar constructions by swapping vocabulary items in different slots of the Cxn. The answer should be the opposite of Base.131313This test is possible for ”Let-alone” and ”Comparative-Correlative”, but is excluded for ”Way-Manner”
4.
Name Bias - this test simply changes the nouns that are referenced in the context of the construction, to see if the model is biased towards interpreting certain Cxns differently depending on the nouns that fill their slots. The answer should be the same as Base.
5.
Short Calibration - this test removes the Cxn altogether, making it impossible to answer the question. The model should respond that it does not have enough information to answer.
6.
Name Calibration - this test gives contextual information unrelated to the construction, in the form of nouns unrelated to the question. The model should respond that it does not have enough information to answer.
7.
Adjective Calibration - this test again provides irrelevant contextual information, in the form of an irrelevant adjective. The model should respond that it does not have enough information to answer.
8.
Nonce Calibration - this test provides irrelevant contextual information in the form of a nonce word instead of meaningful contextual information. This test is the clearest example of a case where the model has no prior reason to provide any answer, and the model should respond that it does not have enough information.
Examples for each of the reasoning tests for each of our constructions can be found below in Table 20.
| Cxn Name | Test Type | Test Text | Answer |
| Comparative Correlative | Base | The more they work, the more I will pay them. The less they work, the less I will pay them. They work more in December than they do in September. How much will I pay them in December, compared to in September? | pay more |
[…]
| Let Alone | Base | Beecher’s reputation as a preacher, let alone as a Man of God, was not universally accepted. Beecher’s reputation as a preacher was weaker in July than in December. When was Beecher more likely to be seen as a Man of God, December or July? | December |
[…]
| Way-Manner | Base | Thus a miner who slipped from the worn and greasy rungs was liable to fall a long way and a dropped tool could clatter its way down to cause death or injury to some fellow below. A dropped tool did more clattering on Monday than on Saturday. When was the dropped tool likely to be falling, on Monday or on Saturday? | Monday |
[…]
Appendix H Annotator Information
All datasets are annotated by co-authors of this paper. 2 annotators identify as men, 4 annotators identify as women. Of our annotators, 3 have a graduate degree in linguistics while 3 do not. At least one linguistic expert and one non-expert annotate each dataset. Because the source of our datasets is web corpora, there is some risk of offensive or hateful content. During annotation, annotators were asked to remove any hateful or offensive content as well as personal identifying information. The annotation task was given to annotators in a spreadsheet. The instructions were identical to model prompts, to ensure maximum comparability to models. All annotators were generally familiar with the tasks of NLI and QA in general, so no general task-related training was needed.
Appendix I Model Hyperparameters
We use the default hyperparameters for both GPT-4o and GPT-3.5, though we set the temperature of model responses to 0, to ensure maximal reproducibility of our work. All our experiments were done through the OpenAI API. The total cost of our experiments was approximately $250 USD.
Appendix J Llama Results
In addition to experiments with GPT-3.5 and GPT-4o we also ran additional experiments using LLaMA-3-8b-instruct and LLaMA-3-70b-instruct and have shared the results in the tables below. Overall, we observe that the trend of this dataset being difficult for LLMs holds.
The patterns we see Table 21 are similar to what we have seen with GPT:
Adding more in-context examples of construction improves performance.
Increasing the number of SNLI examples does not improve performance.
This reinforces our first conclusion that our dataset tests a different axis of semantic knowledge than more general datasets.
In Tables 2, 3 and 4 we see that the performance degrades compared to the performance on the first experiment, similar to the GPT models in the paper. This demonstrates that these tasks are also a difficult problem for LLaMA 3 and therefore act as illustrative examples of LLM capabilities.
In some cases we observe a slight difference in which Cxn/setting each model performs better in. We believe this is a reflection of their instruction fine-tuning data, which is likely to be different given that they have different creators. However, overall we observe that model performance improves with increasing parameter count and we see similar patterns in task difficulty for GPT and LLaMA models which are consistent with the findings in our paper.
| Setting | IC Data | Accuracy | |
| meta-llama-3-8b-instruct | meta-llama-3-70b-instruct | ||
| Zero-shot | None | 0.59 | 0.74 |
| One-shot | CxNLI | 0.65 | 0.84 |
| Three-shot | CxNLI | 0.73 | 0.91 |
| One-shot | SNLI | 0.59 | 0.75 |
| Three-shot | SNLI | 0.58 | 0.74 |
| Setting | IC Data | Accuracy | |
| meta-llama-3-8b-instruct | meta-llama-3-70b-instruct | ||
| Zero-shot | None | 0.59 | 0.74 |
| One-shot | CxNLI-challenge | 0.53 | 0.71 |
| One-shot | CxNLI-challenge, CxNLI | 0.66 | 0.84 |
| Setting | IC Data | Accuracy | |
| meta-llama-3-8b-instruct | meta-llama-3-70b-instruct | ||
| Zero-shot | None | 0.38 | 0.52 |
| One-shot | CxNLI | 0.29 | 0.60 |
| Three-shot | CxNLI | 0.35 | 0.50 |
| One-shot | SNLI | 0.37 | 0.54 |
| Three-shot | SNLI | 0.39 | 0.57 |
| Cxn | Accuracy | |
| meta-llama-3-8b-instruct | meta-llama-3-70b-instruct | |
| Comparative-Correlative | 0.27 | 0.5 |
| Let-Alone | 0.42 | 0.53 |
| Way-Manner | 0.45 | 0.44 |